# PrivTuner with Homomorphic Encryption and LoRA: A P3EFT Scheme for Privacy-Preserving Parameter-Efficient Fine-Tuning of AI Foundation Models

### v1

@@ IX Conclusion
$\mathcal{L}_{2}(\bm{p},\bm{B},\beta,\bm{\gamma})$
$=\sum_{n=1}^{N}\big{(}(p_{n}d_{n})^{2}z_{n}+\frac{1}{4(r_{n})^{2}z_{n}}\big{)}$
$+\beta(\sum_{n=1}^{N}B_{n}-B^{\textnormal{total}})+\sum_{n=1}^{N}\gamma_{n}(r_{n}^{\textnormal{min}}-r_{n}).$ (48)
@@
    $r_{n}^{\textnormal{min}}>\hat{r}_{n}$. Here, we set $\gamma_{n}>0$ and make sure $r_{n}=r_{n}^{\textnormal{min}}$ according to (50b). Given this condition, and by utilizing $\overline{B}_{n}(\beta,\gamma_{n})$ alongside $\overline{p}_{n}(\overline{B}_{n},\gamma_{n})$, we could derive the solution $\gamma_{n}$ denoted as a function of employed $\beta:\gamma_{n}(\beta)$.

### v3

@@ Appendix D Proof of Theorem 1
$\mathcal{L}_{2}(\bm{p},\bm{B},\beta,\bm{\gamma})$ $=\sum_{n=1}^{N}\big((p_{n}d_{n})^{2}z_{n}+\frac{1}{4(r_{n})^{2}z_{n}}\big)$
$\hskip-10.0pt+\beta(\sum_{n=1}^{N}B_{n}-B^{\textnormal{total}})+\sum_{n=1}^{N}\gamma_{n}(r_{n}^{\textnormal{min}}-r_{n}).$ (56)
@@
    $r_{n}^{\textnormal{min}}>\hat{r}_{n}$. Here, we set $\gamma_{n}>0$ and make sure $r_{n}=r_{n}^{\textnormal{min}}$ according to (58b). Given this condition, and by utilizing $\overline{B}_{n}(\beta,\gamma_{n})$ alongside $\overline{p}_{n}(\overline{B}_{n},\gamma_{n})$, we could derive the solution $\gamma_{n}$ denoted as a function of employed $\beta\mathrel{\mathop{\ordinarycolon}}\gamma_{n}(\beta)$.